I do not know how to solve a negative absolute value equation and am looking for help on this such equation. Any help would be greatly appreciated.
-|6-9a|=16

Answer by jim_thompson5910(35256)

-|6-9a|=16 is the same as |6-9a| = -16 after you multiply both sides by -1


Absolute value represents distance the number is from 0.


Distance can never be negative because negative distance makes no sense.


That means |6-9a| = -16 makes no sense. The expression |6-9a| is NEVER negative.


So the equation |6-9a| = -16 has NO solutions.


Consequently, -|6-9a|=16 has NO solutions. No matter what you plug in for 'a', you will never be able to make that equation true.